Eight years ago,Manny was three times as old as Ronnie.Now he is only twice as old as Ronnie.Find their ages.
Found 2 solutions by advanced_Learner, addingup:
Answer by advanced_Learner(501)

now
x ronnie
2x manny
8 years ago
x-8 ronnie
2x-8=3x-24
x=16 and 32

Answer by addingup(3677)

M-8 = 3(R-8) (1)
and
M = 2R Substitute in (1)
2R-8 = 3R-24 Add 24, both sides. And subtract 2R, both sides:
16 = R Ronnie is 16 and Manny is 2(16)= 32